Corp.'s expected year-end dividend is D1 = $1.50, its required return is rS = 12.00%, its dividend yield is 8.00%, and its growth rate is expected to be constant in the future. What is Sorenson's expected stock price in 7 years, i.e., what is P7?
Answer #1

Please find the answer as follows:

D1 = 1.50

Required rate of return (rs) = 12.0%

Dividend Yield = D1/P0 = 8.0%

Growth rate: g = rs - yield = 12 - 8 = 4%

P0 = D1/(rs - g) = 1.5/(.12 - .04) = 18.75

P7 = P0*(1+g)^7 = 18.75*(1+.04)^7= $24.67